OpenStudy (anonymous):

How many grams of nitrogen dioxide may theoretically be produced when copper reacts completely with 21g of nitric acid? Cu + 4HNO3 ---> Cu(NO3)2 + 2H2O + 2NO2

OpenStudy (anonymous):

It says the answer is 7.7 g. I don't know how to get it though. Please guide me through.

OpenStudy (chmvijay):

see 4 moles of HNO3 reacting with Cu produces 2 moles of NO2 right

OpenStudy (anonymous):

ya

OpenStudy (chmvijay):

can you tell molecular mass of HNO3 and NO2

OpenStudy (anonymous):

63 and 46

OpenStudy (anonymous):

So there is 1/3 mol of HN03

OpenStudy (chmvijay):

good :) now i convert those moles to gram using equation moles * molar mass =gram that is 4*63 gram of HNO3 will give rise to 2*46 g of NO2 then 21 gram of HNO3 gives = 21*2*46 /4*63 = 7.66

OpenStudy (anonymous):

that is 4*63g/mol gives rise to 2*46g/mol right?

OpenStudy (anonymous):

@chmvijay

OpenStudy (chmvijay):

Yaaa that also u can doo :) since they have asked in grams i have initially only converted to gram u may also convert moles to gram also :)

OpenStudy (anonymous):

So basically, here we have 4 mol * 63g/mol gives rise to 2 mol * 46g/mol. The mols cancel out and u see that 252 g give rise to 92 g of NO2. Since there is only 21 g of HNO3 then 252/x = 21 implies x = 12. Dividing 92/12 = 7.66666. Correct? @chmvijay

OpenStudy (chmvijay):

yaaa :)